### v1

@@ Abstract
Recent work has established an alternative to traditional multi-layer perceptron neural networks in the form of Kolmogorov-Arnold Networks (KAN). The general KAN framework uses learnable activation functions on the edges of the computational graph followed by summation on nodes. The learnable edge activation functions in the original implementation are basis spline functions (B-Spline). Here, we present a model in which learnable grids of B-Spline activation functions can be replaced by grids of re-weighted sine functions. We show that this leads to better or comparable numerical performance to B-Spline KAN models on the MNIST benchmark, while also providing a substantial speed increase on the order of 4-9 times.
@@ 1 Introduction
In Section section 2 we describe the SineKAN architecture and outline a weight initialization strategy that scales consistently with differing grid sizes and stabilizes numerical performance across multi-layer models. In Section section 3, we present results of model inference speed and performance on the MNIST benchmark and compare it with B-SplineKAN implementation. We discuss our results in Section section 4 and summarize our findings in Section section 5.
@@ 2.1 Sinusoidal Activation Function
where are the layer output features, are the layer input features, is a grid phase shift, are the Fourier coefficient weights, and is a bias term. The base functional form of the sines are fixed, while the functional form of the sine activations is learned through phase and frequency transformations performed by the weight matrix.
@@ 2.2 Grid Phase Shift
Here there is an additional weight matrix compared to our solution. By introducing a phase shift over the grid and input dimensions, we reduce the number of learnable parameters while still allowing for full expression of a Fourier series:
@@ 3.1 Numerical Performance on MNIST
We train and compare single-layer B-SplineKAN and SineKAN networks on the MNIST dataset [29]. We use a batch size 128, learning and decay rates of 0.9 and learning rates of 5e-3 for B-Spline and 3e-4 for SineKAN, optimized with grid search. We test model performance with single hidden layer dimensions of 16, 32, 64, 128, 256, 512 training for 30 epochs using cross entropy loss [30].
@@
| SineKAN | 0.9587 | 0.9718 | 0.9788 | 0.9831 | 0.9853 | 0.9855 |
Fig. 6 shows the model validation accuracy as a function of the number of epochs. The best accuracies are shown in Table 1. The SineKAN model generally achieves comparable results with the B-SplineKAN model for all hidden layer sizes.
We additionally explore fitting using the same hyperparameters but with 1, 2, 3, and 4 hidden layers of size 128. Figure 7 shows that the SineKAN outperforms the B-SplineKAN for all layer depths. Additionally, the SineKAN model accuracy improves at depth 2, 3, and 4 with while the B-SplineKAN validation accuracy is generally worse at higher depths with only comparable performance at depth 3.
An intuitive way to understand the improvement at deep models in the SineKAN framework is through the functional form of a sine expression:
In our model there is no learned frequency term, only a learned amplitude and an on-edge bias. This means the functional form of the first layer simplifies to:
However, at a depth 2 this functional form becomes:
It can also be shown that the functional form of:
where is an amplitude term in the first layer, can also be approximated by:
where and are non-linear functions of . For the simple cases of and , values of and give a reasonable approximation. Therefore, the total functional form at layers beyond the first layer can be approximated as scaled sums of sine functions of the outputs of previous layer.
This functional form for deeper layers then includes the amplitude term, A, from the first layer, as an additional frequency-like term, , in the deep layers. This adds an additional degree of freedom in the model’s approximation of functional mappings that is present in the second layer and beyond but not present in the first layer. This additional degree of freedom contributes to consistently better numerical performance of the model at depth two and beyond.
@@ 3.2 Model Inference Speeds
As Figure 8 shows, the SineKAN has better inference times at all batch sizes and scales better with larger batch sizes, ranging from 4.66 times faster at a batch size of 16 to 9.66 times faster a batch size of 512. While substantially faster at all batch sizes explored, the SineKAN model scales slightly worse with hidden layer size, ranging from around 9.17 at a batch size of 16 to 6.10 times faster at a batch size of 512. Finally, the SineKAN scales slightly better with number of layers, at 7.11 times faster for a depth of one hidden layer and 8.18 times faster for a depth of four hidden layers.
@@ 4 Discussion
In Section section 3.1, we show that our model increasingly outperforms the B-SplineKAN model at very large hidden layer dimensions. We further find that our model generally improves at higher model depths, while B-SplineKAN does not. These results suggest that the SineKAN model may be a better choice compared to B-SplineKAN for scalable, high-depth, large batch models such as large language models (LLMs).
We show in Section section 3.2 that the SineKAN model is on the order of 4 to 9 times faster than an efficient implementation of B-SplineKAN [12] under similar circumstances with better scaling for deep models over large batch sizes. However, while still substantially faster at all hidden layer sizes explored, the scaling is slightly worse than B-SplineKAN for large hidden layer sizes. This can largely be attributed to large memory overhead in the B-SplineKAN operations even at small hidden layer sizes, however, this is worth future exploration.
We also found that SineKAN had a significantly different optimal learning rate compared to B-SplineKAN which motivates the idea that a fair comparison cannot be done across different KAN implementations without performing a grid search to find optimal hyperparameters. Further, we also showed in Section section 2.3, that B-SplineKAN has an inherent flaw in scaling to multi-layer models in that it increasingly constricts layer outputs at higher depths. This likely necessitates additional layer normalization between B-SplineKAN layers. We recommend that approaches for stabilizing model weights at different sizes and depths, similar to those outlined in section 2.3 and section 2.2, should be employed in other KAN models to improve deep model stability and performance.
In summary, sinusoidal activation functions appear to be a promising candidate in the development of Kolmogorov-Arnold models. SineKAN has superior performance in inference speed and accuracy, as well as multi-layer scaling when compared with B-SplineKAN. However, a number of other activation functions mentioned in Section section 1 have also shown to have superior speed and scaling at comparable numerical performance. Further exploration is needed to compare the performance both in terms of inference speed and numerical performance on the broad range of KAN implementations, and on a broader range of tasks, under fair conditions.
@@ 5 Conclusion
We present the SineKAN model, a sinusoidal activation function alternative to B-SplineKAN Kolmogorov-Arnold Networks. We find that this model leads to comparable or better numerical performance on the MNIST benchmark task, when both models are trained used near optimal hyperparameters found with grid search. The SineKAN model increasingly outperforms B-SplineKAN at higher depths and higher hidden dimension sizes. We further find that our model outperforms efficient implementations of B-SplineKAN on speed benchmarks by a factor of four to nine times with better scaling across batch size.

### v2

@@ Abstract
Recent work has established an alternative to traditional multi-layer perceptron neural networks in the form of Kolmogorov-Arnold Networks (KAN). The general KAN framework uses learnable activation functions on the edges of the computational graph followed by summation on nodes. The learnable edge activation functions in the original implementation are basis spline functions (B-Spline). Here, we present a model in which learnable grids of B-Spline activation functions are replaced by grids of re-weighted sine functions. We show that this leads to better or comparable numerical performance to B-Spline KAN models on the MNIST benchmark, while also providing a substantial speed increase on the order of 4-8 times.
@@ 1 Introduction
In Section section 2 we describe the SineKAN architecture, whether it satisfies a universal approximation, and outline a weight initialization strategy that scales consistently with differing grid sizes and stabilizes numerical performance across multi-layer models. In Section section 3, we present results of model inference speed and performance on the MNIST benchmark and compare it with B-SplineKAN implementation. We discuss our results in Section section 4 and summarize our findings in Section section 5.
@@ 2.1 Sinusoidal Activation Function
where are the layer output features, are the layer input features, is a phase shift over the grid and input dimensions, is a grid frequency, are the amplitude weights, and is a bias term. The base functional form of the sines are fixed, while the functional form of the sine activations is learned through learnable frequency and amplitude terms performed over a grid of fixed phases.
@@ 2.2 Grid Phase Shift
Here there is an additional weight matrix compared to our solution:
By introducing learnable frequencies over a grid with fixed phase shifts and input dimensions, we reduce the number of learnable parameters from to where o is the output dimension, i is in the input dimension, and g is the grid size. We show later in section 2.4 that this still satisfies universal approximation.
@@ 2.3 Scaling of Phase Terms With Grid Size
2.4 Universal Approximation
The combination of learnable amplitudes and sinusoidal activation functions have previously been explored [25, 27, 28, 26]. However, these models only satisfy universal approximation on a single layer when combined with linear transformations in sufficiently wide or deep models. By introducing an additional degree of freedom in the form of learnable frequencies over phase-shifted grids, one can eliminate the linear transformation layers.
Any sufficiently well-behaved/smooth 1 dimensional function can be expressed in terms of a Fourier transform w.r.t. a continuous phase space of frequencies :
where and are real-valued functions. The above integral can be discretized using Riemannian sums over a finite set of frequencies where cardinality of the set is the grid size. We henceforth propose the following variational function as an ansatz for :
where we make the replacements , and . Here, are G fixed, finite points from (0,] while we treat all other subscripted symbols as weights whose values can trained to optimize a loss function between and , which converges to the Fourier transform integral of as . Hence, in the limit where , it’s a valid candidate for a learnable activation function ansatz to be used in a Kolmogorov-Arnold network (KAN).
In 6 we show that with a layer with grid size of 100, a single-input, single-output SineKAN layer can map inputs to outputs in 1D functions including over non-smooth functional mappings.
@@ 3.1 Numerical Performance on MNIST
We train and compare single-layer B-SplineKAN and SineKAN networks on the MNIST dataset [29]. We use a batch size 128, learning and decay rates of 0.9 and learning rates of 5e-3 for B-Spline and 4e-4 for SineKAN, optimized with grid search. The models are trained using the AdamW optimizer with a weight decay of 0.01 for B-SplineKAN and 0.5 for SineKAN also found via grid search [30]. We test model performance with single hidden layer dimensions of 16, 32, 64, 128, 256, 512 training for 30 epochs using cross entropy loss [31].
@@
| SineKAN | 0.9618 | 0.9767 | 0.9819 | 0.9851 | 0.9854 | 0.9860 |
Fig. 7 shows the model validation accuracy as a function of the number of epochs. The best accuracies are shown in Table 1. The SineKAN model achieves better results than the B-SplineKAN model for all hidden layer sizes.
We additionally explore fitting using the same hyperparameters but with 1, 2, 3, and 4 hidden layers of size 128. Figure 8 shows that the SineKAN outperforms the B-SplineKAN at lower layer depths. Additionally, the SineKAN model accuracy predictably decreases at higher depths with same hyperparameters while the B-SplineKAN validation accuracy erratically varies at higher depths.
@@ 3.2 Model Inference Speeds
As Figure 9 shows, the SineKAN has better inference times at all batch sizes and scales better with larger batch sizes, ranging from 4.29 times faster at a batch size of 16 to 8.10 times faster a batch size of 512. While substantially faster at all batch sizes explored, the SineKAN model scales slightly worse with hidden layer size, ranging from around 8.09 at a layer size of 16 to 5.73 times faster at a layer size of 512. Finally, the SineKAN scales slightly better with number of layers, at 6.49 times faster for a depth of one hidden layer and 7.14 times faster for a depth of four hidden layers.
@@ 4 Discussion
In Section section 3.1, we show that our model increasingly outperforms the B-SplineKAN model at very large hidden layer dimensions. These results suggest that the SineKAN model may be a better choice compared to B-SplineKAN for scalable, high-depth, large batch models such as large language models (LLMs).
We show in Section section 3.2 that the SineKAN model is on the order of 4 to 8 times faster than an efficient implementation of B-SplineKAN [12] under similar circumstances with better scaling for deep models over large batch sizes. However, while still substantially faster at all hidden layer sizes explored, the scaling is slightly worse than B-SplineKAN for large hidden layer sizes. This can largely be attributed to large memory overhead in the B-SplineKAN operations even at small hidden layer sizes, however, this is worth future exploration.
We also found that SineKAN had a significantly different optimal learning rate and weight decay compared to B-SplineKAN which motivates the idea that a fair comparison cannot be done across different KAN implementations without performing a grid search to find optimal hyperparameters. Further, we also showed in Section section 2.3, that B-SplineKAN has an inherent flaw in scaling to multi-layer models in that it increasingly constricts layer outputs at higher depths. This likely necessitates additional layer normalization between B-SplineKAN layers. We recommend that approaches for stabilizing model weights at different sizes and depths, similar to those outlined in section 2.3 and section 2.2, should be employed in other KAN models to improve deep model stability and performance.
In summary, sinusoidal activation functions appear to be a promising candidate in the development of Kolmogorov-Arnold models. SineKAN has superior performance in inference speed and accuracy, as well as multi-layer scaling when compared with B-SplineKAN. However, a number of other activation functions mentioned in Section section 1 have also shown to have superior inference speed and better numerical performance. Further exploration is needed to compare the performance both in terms of inference speed and numerical performance on the broad range of KAN implementations, and on a broader range of tasks, under fair conditions.
@@ 5 Conclusion
We present the SineKAN model, a sinusoidal activation function alternative to B-SplineKAN Kolmogorov-Arnold Networks. We find that this model leads to better numerical performance on the MNIST benchmark task, when both models are trained using near-optimal hyperparameters found with a parameter grid search. The SineKAN model outperforms B-SplineKAN at higher hidden dimension sizes with more predictable performance scaling at higher depths. We further find that our model outperforms efficient implementations of B-SplineKAN on speed benchmarks by a factor of four to nine times with better scaling across batch size.